# An Empirical Study of Developer-Provided Context for AI Coding Assistants in Open-Source Projects

### v1

@@ 5.3. RQ3: Context Co-occurrence
At a more granular level, developers usually provide a selected amount of codes in their cursor rules to prevent overwhelming LLMs with excessive context and keep LLM focus on the most critical aspects of the project. Figure 7 shows the distribution between the number of codes covered by a single repository. Very few repositories provide more than 19 unique codes and most of the repositories (323 repositories, 80.55%) provide a moderate amount of unique code types (between 3 and 15 types) with a peak at 9 unique code types. This indicates that developers tend to provide a balanced amount of context, covering multiple aspects of the project. At the same time, this also suggests that developers are selective in the specific codes they include, or they miss out on certain aspects that could be beneficial for LLMs to understand the project better, which aligns with prior findings on documentation practices (maalej2013patterns).
5.4. RQ4: Proportion of Duplicated Lines
Figure 8. Percentage of code that are duplicated by context category.
Findings. Overall, we identified a total of 19917 (28.70%) duplicated lines across all repositories with an average of 49.67 duplicate lines (SD=164.20) per repository. Figure 8 shows the percentage of code that are duplicated by context category.
Figure 9 shows the distribution of the average number of duplicated lines for each context type by programming language. We found that different programming languages tend to have different copying practices when writing cursor rules. For example, developers using programming languages for desktop and enterprise applications, such as C#, which require high domain specificity and are less tolerant of errors, tend to replicate cursor rules from similar projects more frequently (M=61.14, SD=161.77). In contrast, developers using languages for frontend applications, such as TypeScript, often copy bits and pieces from various sources but also contribute a large amount of original content, resulting in the highest total number of duplicated lines (7,838 lines).
For individual repositories that are not complete duplicates of others, we found an average duplication rate of 16.30% of lines, with the top 10 repositories having over 96% of their lines duplicated. This suggests that some developers tend to extensively reuse existing cursor rules when writing their own, rather than creating everything from scratch.
@@ 5.5. RQ5: Cursor Rule Composition Over Time
Projects created two to five years before the introduction of cursor rules tend to provide the largest amount of guideline-related context. This may be because these projects are old enough to require more context to ensure LLMs understand legacy constraints, but not so old that their technical guidelines have already become widely established online.
@@
Limitations. According to official guidance for writing cursor rules, developers are required to place their cursor rules in the .cursor/ directory by default for Cursor to index. Therefore, we only checked the .cursor/ folder, which may miss monorepos or custom structures. This bias is consistent across all groups and does not affect the relative comparisons.
@@ 6. Discussion and Implications
6.1. Rule Authoring
The Need for Broader Context Support. Current AI-powered code editors typically utilize context retrieved from the IDE or from user-highlighted code. Our data show that developers actively provide additional context beyond source code, such as conventions (15%), guidelines (33%), and even instructions on how the LLM should behave (8%). This aligns with the emerging trend in popular AI assistants to expand their context mechanisms to ingest non-code artifacts—such as documentation, configuration files, and explicit rule files—to improve the relevance and accuracy of generated responses.
@@ 6.2. Tool Building
Context-Aware Rule Generation. We observed that the types of rules that are most effective can vary by programming language or application domain. For example, statically typed languages like C# tend to require more project-specific context, while frontend languages like JavaScript/TypeScript benefit from more examples. However, developers are currently adopting existing rule templates (28.7% of all lines) or using LLMs to help generate their cursor rules, as observed in at least one repository among the 30 we initially coded. Cursor also officially provides a command to help developers generate initial cursor rules using LLMs. Without proper context awareness, these generated rules may not be optimal for the specific characteristics of each project. This indicates significant potential for future tools to generate more context-aware rules, tailored to the specific characteristics of each project, rather than relying on generic templates. This could further increase the efficiency of the entire development process. Further research on automatically identifying necessary domain- or language-specific context is needed to inform the design of such tools. Future work can also focus on building better cursor rule generation or context assessment tools to help developers create and evaluate the comprehensiveness of their provided context. Additionally, computer use agents could further support agentic systems by automatically generating context rules based on project structure and existing documentation.
6.3. Future Research Directions
@@
Qualifying the Efficacy of Context Rules. The rules we observed are primarily based on developer intuition; their actual impact on LLM performance remains an open question. For example, it is unclear whether project-related context is more beneficial than allowing LLMs to automatically index project files. In rare cases, excessive context may even degrade performance if it confuses the LLM. Future work is needed to quantify how different types of context—such as the themes in our taxonomy—contribute to task performance and to determine the optimal amount for each context type. Similar to how prompt engineering techniques are empirically tested, the ”usefulness” of different context types could be evaluated by comparing end-task results with and without certain context types. This would help prioritize which context is most valuable, especially since authoring and maintaining these rules imposes a cognitive load on developers.
@@
Longitudinal Analysis. Our study provides a cross-sectional snapshot of how developers use cursor rules. However, this is a rapidly evolving practice. As developers become more familiar with LLM capabilities, their use of these rules may change. For example, they may shift from providing detailed project information to focusing more on LLM-specific instructions as they become more aware of LLM capabilities. Furthermore, as LLMs themselves become more powerful, the relative importance of different context types may shift. A longitudinal study is needed to track the evolution of these rules over time. One potential approach, common in the MSR community, is to monitor changes in cursor rules as a proxy for their effectiveness. Observing whether repositories abandon, expand, or refine their rules would provide critical insights into their perceived usefulness and the changing nature of human-AI collaboration in software engineering.
@@
Future Education. With the emergence of AI in development workflows across all areas, educating developers on how to effectively collaborate with AI systems is critical. Our analysis of creation time differences also shows that newly created repositories often have less context, suggesting that developers may be unsure about what needs to be included. This highlights the need for further education. Our findings reveal several dimensions of how developers currently provide context to LLMs, which can inform future educational materials instead of treating AI as a magic box. For example, developers may benefit from learning about the importance of balancing different context types, such as project information, conventions, guidelines, and LLM directives. Future educational resources should help developers understand and apply best practices when authoring context for LLM-powered coding assistants.
@@ 7. Conclusion
In this paper, we presented a large-scale empirical study of 401 open-source repositories containing cursor rules to understand what types of context developers provide to LLM-powered coding assistants. We identified five main categories of context: Project Information, Conventions, Guidelines, LLM Directives and Examples, with some overlapping information traditionally shared between human developers and others specific to LLMs. We believe our findings, particularly the detailed taxonomy, provide a valuable foundation for the design of future LLM-powered coding assistants, as well as for understanding the knowledge and information that will be shared within teams as LLMs become new collaborators.

### v2

@@ 5.3. RQ3: Proportion of Duplicated Lines
For individual repositories that are not complete duplicates of others, the average line duplication rate is 16.30%, with the top 10 exceeding 96%. This suggests that some developers extensively reuse existing cursor rules, adding additional context as needed.
@@
5.4. RQ4: Context Co-occurrence
To test whether developers provide broad, multi-aspect context in cursor rules, we analyzed the co-occurrence of context categories within repositories and the number of unique codes per repository.
Findings. Figure 8 illustrates the co-occurrence of context categories in cursor rules. Overall, developers tend to provide comprehensive context for LLMs: 149 repositories (37.16%) include all four core content categories (excluding examples), and 99 repositories (24.69%) include all five categories.
Developers prefer to write cursor rules in a style similar to traditional project or collaboration documentation, which typically covers best practices and standards for software development (prana2019categorizing; fronchetti2023contributing). In our case, this includes project, conventions, and guidelines category. People usually pay the same amount of attention to the tuple combinations of those three context types, which were relatively evenly distributed (Project and Convention: 296, 73.8%; Project and Guideline: 310, 77.3%; Guideline and Convention: 311, 77.6%). Examples are usually provided by the traditional documentation to illustrate other context types and help people use the other context types more effectively as well (exampleeffectiveness). In our case, the Example category is usually co-occurs with all other three categories: Guideline, LLM Directive, and Convention (172, 42.9%). Individually, examples are most frequently provided alongside Guidelines (193, 48.1%), Conventions (189, 47.1%), LLM Directives (187, 46.6%)
Some developers choose to only focus on specific aspects when writing cursor rules to let LLM focus on a really specific aspects while generating code. 32 repositories (7.98%) provided only one category of context in their cursor rules; most of them were for guideline-related context (11, 2.74%).
At a more granular level, developers typically provide a selective number of codes in their cursor rules to avoid overwhelming LLMs and to keep the model focused on the most critical aspects of the project. Figure 8 shows the distribution of the number of unique codes per repository. Very few repositories provide more than 19 unique codes, and most repositories (323, 80.55%) include a moderate number of unique codes (between 3 and 15), with a peak at 9 unique codes. This suggests that developers balance context coverage across the project but are selective in their choices, which is consistent with prior findings on documentation practices (maalej2013patterns).
@@ 5.5. RQ5: Cursor Rule Composition Over Time
Limitations. We only checked the .cursor/ folder following official guidance. This may miss monorepos or custom structures, but the bias is consistent and does not affect comparisons.
@@ 6. Discussion and Implications
Context-Aware Rule Generation. We observed that the types of rules that developers think are most effective can vary by programming language or application domain. For example, cursor rules for statically typed languages like C# tend to have more project-specific context, while rules for frontend languages like JavaScript/TypeScript had more examples. However, developers are currently adopting existing rule templates (28.7% of all lines) or using LLMs to help generate their cursor rules. Cursor also officially provides a command to help developers generate initial cursor rules using LLMs. Without proper context awareness, these generated rules may not be optimal for the specific characteristics of each project. This indicates significant potential for future tools to generate more context-aware rules, tailored to the specific characteristics of each project, rather than relying on generic templates. This could further increase the efficiency of the entire development process. Further research on automatically identifying necessary domain- or language-specific context is needed to inform the design of such tools. Future work can also focus on building better cursor rule generation or context assessment tools to help developers create and evaluate the comprehensiveness of their provided context.
@@
Qualifying the Efficacy of Context Rules. The rules we observed are primarily based on developer intuition; their actual impact on LLM performance remains an open question. For example, it is unclear whether project-related context is more beneficial than allowing LLMs to automatically index project files, as excessive context may degrade LLM performance. Future work is needed to quantify how different types of context—such as the themes in our taxonomy—contribute to task performance and to determine the optimal amount for each context type. Similar to how prompt engineering techniques are empirically tested, the ”usefulness” of different context types could be evaluated by comparing end-task results with and without certain context types. This would help prioritize which context is most valuable, especially since authoring and maintaining these rules imposes a cognitive load on developers.
@@
Longitudinal Analysis. Our study provides a cross-sectional snapshot of how developers use cursor rules. However, this is a rapidly evolving practice. As developers become more familiar with LLM capabilities, their use of these rules may change. For example, they may shift from providing detailed project information to focusing more on LLM-specific instructions. Furthermore, as LLMs are upgraded, the relative importance of different context types may shift. A longitudinal study is needed to track the evolution of these rules. One potential approach, common in the MSR community, is to observe whether repositories abandon, expand, or refine their rules as a proxy for the perceived usefulness and the changing nature of human-AI collaboration in software engineering.
@@
Future Education. Our analysis of creation-time differences shows that newly created repositories often contain less context, suggesting that developers are uncertain about what is necessary. This highlights the need for further education. Our findings across multiple dimensions of developer-provided context can inform educational materials, which should help developers better understand and apply best practices.
@@ 7. Conclusion
In this paper, we presented a large-scale empirical study of 401 open-source repositories containing cursor rules to understand what types of context developers provide to LLM-powered coding assistants. We identified five main categories of context: Project Information, Convention, Guideline, LLM Directive and Example, with some overlapping information traditionally shared between human developers and others specific to LLMs. We believe our findings, particularly the detailed taxonomy, provide a valuable foundation for the design of future LLM-powered coding assistants, as well as for understanding the knowledge and information that will be shared within teams as LLMs become new collaborators.
Data Availability. Our supplementary material, the dataset and the prompts, is available at DOI https://doi.org/10.5281/zenodo.18313203  (jiang_2026_18313203).